OpenStudy (anonymous):

in a class with a certain number of students , if one student weighing 50kg is added then average weight of class incereases by 1 kg ,if one more student weighin 50kg is added then the average weight of class increasses by 1.5kg over the original average? what isthe original average weight (in kg) of the class?

OpenStudy (anonymous):

option are.....4 46 2 47

OpenStudy (anonymous):

a.4 b.46 c.2 d.47

OpenStudy (anonymous):

we have ( Xi + 50 ) / ( n+1) = Xi / n

OpenStudy (anonymous):

( Xi + 50 ) / ( n+1) = Xi / n + 1

OpenStudy (anonymous):

how to do

OpenStudy (anonymous):

its a long one, there might be a trick

OpenStudy (anonymous):

15. Ralph likes 25 but not 24; he likes 400 but not 300; he likes 144 but not 145. Which does he like: 10 50 124 200 1600

OpenStudy (anonymous):

well what abt......... ‎976 : 684 375 : 225 864 : 512 597 : ?

OpenStudy (anonymous):

you have 2 equations with 2 variables : the initial average weight (Xi) and the initial number of people (n). The 2 equations are : (Xi + 50) / (n+1) = Xi + 1 (Xi + 100) / (n+2) = Xi + 1.5 I'd isolate n and solve it from there

OpenStudy (anonymous):

yea i got answer

OpenStudy (anonymous):

its 47.